• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 19:32:19 on June21,2002

    Entry number 86544

    keyword=cryotarget cool-donw completed

    Cryotarget cool-down completed at 19:30. Switched to cold-return around 18:30.
    Loop1 is LD2 at 22K, ~27.0 psi. Since gas vented soem during warm up, we had to add some. The D2 gas bottle is empty now. Loop1 fan is set at 60 Hz.
    Loop 2 is LH2 at 19K, ~25.8 psi. Fan is 24 Hz.
    Vac gauge tripped off earlier, reset by Gary Dezern. It is reading ~5*10-7.